OpenStudy (anonymous):

Evaluate f(x) = -x^2 + 1 for x = - 3.

OpenStudy (anonymous):

i have no idea what to do

OpenStudy (mokeira):

you already have the value of x...substitute \[-(-3^{2})+1=f(x)\] \[-(9)+1=f(x)\] that should be easier...now just finish it up

OpenStudy (mokeira):

i forgot to substitute the x in f(x)...remember that one!

OpenStudy (jmark):

Given f(x)=-x^2+1 for x=-3 f(-3)=-(-3)^2+1 f(-3)=-9+1 f(-3)=-8 CBSE Computer Science Syllabus for Class 12 - http://scoremore.beep.com/cbse-computer-science-syllabus-for-class-12-2014-08-08-1.htm